OpenStudy (anonymous):

Taylor's doctor has put him on a diet of 800 calories per meal. This means he can consume no more than 800 calories at any meal. His options for lunch in the cafeteria are as follows: Soups that have 150 calories per serving Mini Burgers that have 100 calories per serving Which inequality below can be used to find the possible combinations of servings of soup (s) and mini burgers (b) he can eat and stay within his goal of no more than 800 calories? 150s + 100b 800 150s + 100b 800 150s − 100b 800 150s − 100b 800

OpenStudy (ganpat):

150s + 100b 800 150s + 100b 800 150s − 100b 800 150s − 100b 800 little incomplete but ok :)

OpenStudy (ganpat):

150s + 100b <= 800

OpenStudy (anonymous):

\[150s + 100b \ge 800\] \[150s+100b \le 800\] \[150s - 100b \le 800\] \[150s - 100b \ge 800\]

OpenStudy (anonymous):

@Ganpat

OpenStudy (anonymous):

i found the error

OpenStudy (ganpat):

ha ha, good

OpenStudy (ganpat):

so he can have, less than 800 calories.. so what do u think ??

OpenStudy (anonymous):

i think it's the first one or the last one

OpenStudy (ganpat):

it will never be (-) as he is eating both.. and calories should not exceed more than 800.. i mean should be <= 800

OpenStudy (anonymous):

so it's the first one

OpenStudy (ganpat):

how about second one ??

OpenStudy (anonymous):

oh

OpenStudy (ganpat):

so can it be ??

OpenStudy (anonymous):

yeah

OpenStudy (ganpat):

cool :)

OpenStudy (anonymous):

thanks

OpenStudy (ganpat):

anytime..